one number is two more than a second number, twice the first is four less than three times the second. find the numbers. i need help. :(
Answer by ewatrrr(10682)

Hi,
Let x be the second number number. then (x+2) would be the first number
.
question states: write as you read
2*(x+2) = 3*x -4
.
simplify and solve
2x + 4 = 3x - 4
8 = x (second number)
.
(x+2) = 10 (fist number)
Numbers are 8 and 10
.
check your answer
2*10 = 3*8 - 4 = 20